the number of number pairs (x,y) which will satisfy the equation x2-xy+y2=4(x+y-4)

Grade:11

1 Answers

Arun
4 years ago
 

x2-xy+y2=4(x+y-4)

  x2-x(y+4)+y-4y +16=0

if it has real solution then D≥0

so (y+4)2 -4( y-4y +16)≥0

 or (y-4)2  ≤0

 only possible when y=4

so for y=4 only one value of x=4 possible

so only one pair (4,4) is the solution

 

Hope it helps

 

Regards

Arun (askIITians forum expert)